Reduce the quadratic form

2 2 2 8 7 3 12 – 8 4 x y z xy yz zx    

to the canonical form

through an orthogonal transformation and hence show that it

is positive Semi-definite.



1
Expert's answer
2021-10-20T17:55:11-0400

"8x^2+7y^2+3z^2-8yz+4xz+12xy"

matrix in quadratic form:

"A=\begin{pmatrix}
 8 & 6&2 \\
 6 & 7&-4 \\
2&-4&3
\end{pmatrix}"


"\begin{vmatrix}
 8 -\lambda& 6&2 \\
 6 & 7-\lambda&-4 \\
2&-4&3-\lambda
\end{vmatrix}=0"


"(8-\lambda)((7-\lambda)(3-\lambda)-16)-6(18-6\lambda+8)+2(-24-14+2\lambda)=0"

"(8-\lambda)(21-10\lambda+\lambda^2-16)+36\lambda-156+4\lambda-76=0"

"40-80\lambda+8\lambda^2-5\lambda+10\lambda^2-\lambda^3+40\lambda-232=0"

"\lambda^3-18\lambda^2+45\lambda+182=0"

"\lambda_1=-2.1,\lambda_2=6.3,\lambda_3=13.8"


eigenvectors:


for "\lambda_1=-2.1" :

"10.1x+6y+2z=0"

"6x+9.1y-4z=0"

"2x-4y+5.1z=0"


"26.2x+21.1y=0"

"21.1y-19.3z=0"

"y=-1.2x,z=1.1y"


"x_1=\begin{pmatrix}
 1 \\
 -1.2 \\
-1.4
\end{pmatrix}"


for "\lambda_2=6.3" :

"1.7x+6y+2z=0"

"6x+0.7y-4z=0"

"2x-4y-3.3z=0"


"2.6x-11.3y=0"

"12.7y+5.9z=0"

"y=0.2x,z=-2.2y"


"x_2=\begin{pmatrix}
 1 \\
 0.2 \\
-0.4
\end{pmatrix}"


for "\lambda_3=13.8" :

"-5.8x+6y+2z=0"

"6x-6.8y-4z=0"

"2x-4y-10.8z=0"


"17.6x-18.8y=0"

"5.2y+28.4z=0"

"y=0.9x,z=-0.2y"


"x_3=\begin{pmatrix}
 1 \\
 0.9 \\
-0.2
\end{pmatrix}"


​normalized matrix:

"N=\begin{pmatrix}
 1/|x_1| & 1/|x_2| &1/|x_3|\\
 -1.2/|x_1| & 0.2/|x_2|&0.9/|x_3|\\
-1.4/|x_1|&-0.4/|x_2|&-0.2/|x_3|
\end{pmatrix}"


"|x_1|=2.1,|x_2|=1.1,|x_3|=1.4"


"N=\begin{pmatrix}
 0.5 & 0.9 &0.7\\
 -0.6 & 0.2&0.6\\
-0.7&-0.4&-0.1
\end{pmatrix}"


"N^T=\begin{pmatrix}
 0.5 & -0.6 &-0.7\\
 0.9 & 0.2&-0.4\\
0.7&0.6&-0.1
\end{pmatrix}"


"AN=\begin{pmatrix}
 8 & 6&2 \\
 6 & 7&-4 \\
2&-4&3
\end{pmatrix}\begin{pmatrix}
 0.5 & 0.9 &0.7\\
 -0.6 & 0.2&0.6\\
-0.7&-0.4&-0.1
\end{pmatrix}=\begin{pmatrix}
 -1 & 7.6&9 \\
 1.6 & 7&8.8 \\
-2.4&-0.2&-1.3
\end{pmatrix}"


"D=N^TAN=\begin{pmatrix}
 0.5 & -0.6 &-0.7\\
 0.9 & 0.2&-0.4\\
0.7&0.6&-0.1
\end{pmatrix}\begin{pmatrix}
 -1 & 7.6&9 \\
 1.6 & 7&8.8 \\
-2.4&-0.2&-1.3
\end{pmatrix}="


"=\begin{pmatrix}
 -8.4 &0 &0\\
 0 & 8.3&0\\
0&0&11.7
\end{pmatrix}"


Canonical form: "0.2x^2+8.3y^2+11.7z^2"


quadratic form Q(x) = xT · A · x is positive semidefinite if Q(x) ≥ 0